# Exposing Lip-syncing Deepfakes from Mouth Inconsistencies

### v1

@@ Abstract
A lip-syncing deepfake is a digitally manipulated video in which a person’s lip movements are created convincingly using AI models to match altered or entirely new audio. Lip-syncing deepfakes are a dangerous type of deepfakes as the artifacts are limited to the lip region and more difficult to discern. In this paper, we describe a novel approach, LIP-syncing detection based on mouth INConsistency (LIPINC), for lip-syncing deepfake detection by identifying temporal inconsistencies in the mouth region. These inconsistencies are seen in the adjacent frames and throughout the video. Our model can successfully capture these irregularities and outperforms the state-of-the-art methods on several benchmark deepfake datasets.
Index Terms—  DeepFake detection, Lip-syncing deepfakes, Spatial-temporal inconsistency
1 Introduction
@@
Deepfakes are digitally altered videos in which a person’s appearance and voice are modified to mimic someone else, often for realistic impersonations, leveraging advanced generative AI methods. Various forms of Deepfakes include face swapping [1], face reenactment [2], lip-syncing[3], and face animation [4]. While there are positive uses of deepfakes, such as in movie dubbing, reviving historical figures, and creating virtual avatars for individuals with speech impairments by synthesizing voices similar to their own, deepfakes pose significant societal risks. They can be misused for identity theft, fabricating fake news, or generating fraudulent videos for purposes like extortion or spreading disinformation. Accessible open-source tools111e.g., DeepFaceLab https://github.com/iperov/DeepFaceLab, Wav2Lip https://github.com/Rudrabha/Wav2Lip, FakeApp https://www.malavida.com/en/soft/fakeapp/. allow the creation of deepfake videos featuring celebrities or public figures with minimal technical expertise. A notable recent instance is the circulation of deepfake videos on social media, falsely depicting CNN and Fox News anchors endorsing novel or simplistic diabetes treatments [^2] https://today.rtl.lu/news/fact-check/a/2136220.html.
@@
Fig. 1: Illustration of the mouth inconsistency in lip-syncing deepfakes. We visualize video frames from FakeAVCeleb [5] dataset with open mouths. The first three columns present consecutive frames for local comparison, while the last two columns offer a broader perspective by displaying frames with similar poses from the entire video, defined as global inconsistencies in our paper. The deepfakes exhibit more pronounced inconsistencies in aspects such as mouth shape, coloration, dental structure, and tongue appearance.
@@
Deepfakes are broadly categorized into two types based on the regions of manipulation: entire-face synthesis and partial manipulation. Entire-face synthesis encompasses techniques such as face-swapping and talking head generation. On the other hand, partial manipulation primarily involves lip-syncing videos. In these videos, the lip movements of an individual are altered to match a specific audio track, creating the illusion that they are speaking words they originally did not. Lip-synced deepfakes focus on modifying only a person’s mouth area, which facilitates the creation of high-quality deepfakes. A recent investigation [6] indicates that lip-synced deepfakes pose a greater challenge for human detection than face-swapping deepfakes. However, unlike the extensively researched face-swap deepfake detection [7, 8, 9, 10, 11, 12], identifying lip-syncing deepfakes presents greater challenges and is less explored.
Fig. 2: End to end pipeline of the proposed LIPINC model.
@@
2 Related Work
DeepFake Detection. Most deepfake detection approaches focus on entire-face synthesis in videos, identifying visual anomalies like face warping [9], facial movements [17], discrepancies between face and context [18], and inconsistencies in face identity [19]. These methods extract distinctive features from the entire face and are less effective when dealing with partially manipulated lip-syncing videos. For lip-sync deepfake detection, various techniques are employed to exploit audio-visual mismatches. Shahzad et al.[20] proposed a method to detect fabricated videos by contrasting the real lip sequence from a video with a synthetic lip sequence created based on audio using a lip-sync model [3]. Feng et al.[14] investigated audio-visual signal inconsistencies in forged videos through anomaly detection, but this approach might miss manipulations that achieve good synchronization [14]. Alternatively, Haliassos et al.[15] identified forged videos by detecting semantic anomalies in mouth movements from video segments, focusing on irregularities in mouth motion from lower face regions, which may be less effective in videos with minimal mouth movement. LipForensics [15], in contrast, learned mouth motion patterns using lipreading networks from lower-face clips for deepfake identification. Tolosana et al.[16] examined visual artifacts across various facial regions (eyes, nose, and mouth), utilizing spatial indicators to differentiate between real and fake videos.
Lip-syncing Deepfake Datasets. Early deepfake datasets are for entire-face synthesis, such as face-swapping and face-reenactment [2, 21]. The emergence of Wav2Lip [3] in 2020 marked a significant advancement in lip-syncing deepfakes, and led to the development of several datasets aimed at detecting partially manipulated lip-syncing videos. For instance, the FakeAVCeleb [5] dataset is a comprehensive audio-video deepfake detection resource featuring accurately lip-synced videos. This dataset encompasses three categories of audio-video deepfakes sourced from English-language YouTube videos, including lip-syncing[3], face-swapping[1, 22], and voice cloning [23]. Additionally, the KODF [24] dataset presents a vast collection of both generated and authentic videos with Korean subjects, with the fake videos produced using Wav2Lip and various face-swapping models. Note that these existing deepfake datasets encompass only a subset of lip-syncing videos created with a single model.
3 Method
@@
The complete workflow of the LIPINC model is depicted in Fig.2. It comprises two primary modules: 1) Local and Global Mouth Frame Extractor, which isolates adjacent and similarly posed mouth segments based on mouth openness throughout the video sequence; and 2) Mouth Spatial-Temporal Inconsistency Extractor (MSTIE), responsible for encoding and learning distinctive inconsistency features within and across frames, focusing on both color and residual aspects. The MSTIE module’s output is then directed to a binary classifier to determine the probabilities of each class.
3.1 Local and Global Mouth Frame Extractor
Given an input video, we first use a face detector, such as Dlib [25], to crop and align the face in each video. Then we use the facial landmarks to extract the mouth region, as shown in Fig. 3. We propose to extract multiple frames with open mouths since they contain more inconsistencies regarding deepfakes. For the adjacent mouth extraction as local inconsistency analysis, we obtain a sequence of $L$ adjacent frames with open mouth, which are selected based on a large distance between landmarks 63 and 67 in Fig. 3. For the global inconsistency representation, we search for $G$ similar pose frames identical to the mouth openness in local frames from the rest of the video. To ensure a global matching and prevent selected frames from being adjacent, we set the extraction of similar-pose frames to have a minimum time gap of 0.09 seconds between them. We then concatenate the adjacent and similar-pose frames to form our color sequence C with $N$ frames where $\textbf{C}\in R^{(N*H*W*3)}$, 3 is the channel number, and $N=L+G$. To emphasize the consistency description at the mouth structure level, we further add the structure branch with residual sequence S, where $\textbf{S}\in R^{((N-1)*H*W*3)}$ is calculated by
$S_{t}=C_{t+1}-C_{t}\;(1\leq t<N)$
The color and structure sequences with local and global mouth frames are fed into the following inconsistency feature extraction module.
@@ 3.1 Local and Global Mouth Frame Extractor
Table 1: In-domain testing on FakeAVCeleb dataset [5] and Cross-domain testing on KoDF [24] and LSR+W2L. We compare our model with state-of-the-art detectors regarding Precision, Accuracy, Average Precision and AUC score. Here RI means Re-Implemented. Note that the results sourced from existing methods used much more training data than us for in-domain cross-type testing.
3.2 Mouth Spatial-Temporal Inconsistency Extractor
This MSTIE module is designed to encode the color and structure frames and learn discriminative features representing spatial and temporal inconsistency levels for deepfake detection. We use the 3D-CNN [28] model to produce the spatial-temporal features $\textbf{v}_{ci}$ and $\textbf{v}_{si}$ for the color and the structure sequence, respectively, where $i$ is the number of input video. These vectors are then passed through a cross-attention module which interconnects the color and structure branches such that $K_{ci}=V_{ci}=\textbf{v}_{ci},$ $Q_{ci}=\textbf{v}_{si}$ and $K_{si}=V_{si}=\textbf{v}_{si}$, $Q_{si}=\textbf{v}_{ci}$, where $K,V$ and $Q$ represent the keys, values, and queries. The cross attention for color $A_{c}$ and structure $A_{s}$ branch is calculated as follows:
$A_{ci}=\text{softmax}\left(\frac{Q_{ci}K_{ci}^{T}}{\sqrt{d_{qc}}}\right)V_{ci},A_{si}=\text{softmax}\left(\frac{Q_{si}K_{si}^{T}}{\sqrt{d_{qs}}}\right)V_{si}$
@@ 3.2 Mouth Spatial-Temporal Inconsistency Extractor
$A_{fi}=\text{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{Q}}}\right)V_{i}+\textbf{v}_{si}$
@@
Fig. 3: Mouth region landmarks detected by Dlib [25]. Orange colors denote the landmarks for mouth openness measurement and matching.
3.3 Loss Function
@@
$L_{IL}=-\frac{1}{M}\sum_{i=1}^{M}y_{i}log(AvgS_{i})+(1-y_{i})log(1-AvgS_{i}),$
@@
4 Experiments
Fig. 4: Ablation analysis on LSR+W2L and KoDF-LS Dataset based on AUC scores.Here L,G,C, and S refer to Local, global, Color, and Structure frames, respectively. MSTIE is the Mouth Spatial-Temporal Inconsistency Extractor module, and IL is inconsistency loss
4.1 Settings
Datasets. We use the FakeAVCeleb dataset[5] to train and validate our model. The FakeAVCeleb dataset consists of 500 real unique frontal face videos and more than 20000 multi-modal deepfake videos. We slit the dataset into two categories: Lip-synced (FakeAV-LS) deepfakes that utilize Wav2lip [3] and Face-swap (FakeAV-FS) deepfakes that have not used Wav2lip. Since our method mainly focuses on lip-syncing deepfakes detection, we have divided the lip-synced deepfakes in this dataset into a non-overlapped training set and a test set. We use two more datasets to evaluate our model’s generalization ability. The first one is the Kodf dataset [24], which includes 62,166 real videos and 175,776 deepfake videos. Similar to the FakeAVCeleb dataset, deepfake videos contain lip-syncing and entire-face synthesis. We created 3 test sets from the dataset based on the deepfake generation method. 1) KODF-LS: audio-driven lip-synced deepfakes from [3, 30]; 2) KODF-FSGAN: Faceswap deepfakes generated using [22]; 3) KODF-DFL:Faceswap deepfakes generated using [31]. We randomly selected a subset of real and fake videos for evaluation following previous studies [14], and use 800 samples for each class. Secondly, we have generated a new dataset called LSR+W2L, which is specifically tailored for lip-syncing detection. We consider two models, Wav2Lip and Wav2Lip with GAN model [3]. This dataset contains 1,000 real videos from the LRS2 dataset [32] and 1,000 fake videos.
Evaluation Metrics. The model is evaluated using three widely used metrics, including Precision, Accuracy, Average Precision (AP), and Area Under the Curve (AUC) scores [33].
@@ 4.2 Results
Comparison under in-domain testing. In Table 1, we first evaluate our model on the FakeAvceleb testing set (FakeAV-LS) [5] and compared its performance with the state-of-the-Art (SOTA) methods. The proposed LIPINC model achieves the best performance in terms of accuracy, precision and is tied with the FTCN [7] model for the best performance in terms of AUC scores. Our model outperforms the best SOTA method Intra-Cross-modal [26] in terms of AUC scores by 1% and performs slightly worse than it in AP. Overall, the LIPINC model demonstrates outstanding performance in detecting lip-syncing videos under in-domain testing, thanks to the utilization of spatial-temporal mouth inconsistency features.
@@
Comparison under cross-domain testing. We further evaluate our model on KODF-LS[24] and LSR+W2L datasets to evaluate the generalization capability of our model. The results are shown in Table 1. Note that our model was trained on the FakeAvceleb training set (FakeAV-LS), while the SOTA results cited from existing papers were trained on the whole FakeAVCeleb dataset [5] as introduced in [14]. On our newly created LSR+W2L dataset, the proposed LIPINC model outperforms all the baselines in terms of AP and AUC scores. The RealForensics model achieves better performance than our method on the KODF dataset due to its utilization of a large training dataset. Our method achieves an AUC of over 87% on both datasets, demonstrating strong generalization in detecting unknown lip-syncing videos.

### v2

@@ Abstract
A lip-syncing deepfake is a digitally manipulated video in which a person’s lip movements are created convincingly using AI models to match altered or entirely new audio. Lip-syncing deepfakes are a dangerous type of deepfakes as the artifacts are limited to the lip region and more difficult to discern. In this paper, we describe a novel approach, LIP-syncing detection based on mouth INConsistency (LIPINC), for lip-syncing deepfake detection by identifying temporal inconsistencies in the mouth region. These inconsistencies are seen in the adjacent frames and throughout the video. Our model can successfully capture these irregularities and outperforms the state-of-the-art methods on several benchmark deepfake datasets. Code is available at https://github.com/skrantidatta/LIPINC.
Index Terms:
DeepFake detection, Lip-syncing deepfakes, Spatial-temporal inconsistency
I Introduction
@@
Deepfakes are digitally altered videos in which a person’s appearance and voice are modified to mimic someone else, often for realistic impersonations, leveraging advanced generative AI methods. Various forms of Deepfakes include face swapping [1], face reenactment[2], lip-syncing[3], and face animation[4]. While there are positive uses of deepfakes, such as in movie dubbing, reviving historical figures, and creating virtual avatars for individuals with speech impairments by synthesizing voices similar to their own, deepfakes pose significant societal risks. They can be misused for identity theft, fabricating fake news, or generating fraudulent videos for purposes like extortion or spreading disinformation. Accessible open-source tools111e.g., Wav2Lip: https://github.com/Rudrabha/Wav2Lip,
DeepFaceLab: https://github.com/iperov/DeepFaceLab,
FakeApp: https://www.malavida.com/en/soft/fakeapp/.
allow the creation of deepfake videos featuring celebrities or public figures with minimal technical expertise. A notable recent instance is the circulation of deepfake videos on social media, falsely depicting CNN and Fox News anchors endorsing novel or simplistic diabetes treatments [^2] https://today.rtl.lu/news/fact-check/a/2136220.html.
@@
Figure 1: Illustration of the mouth inconsistency in lip-syncing deepfakes. We visualize video frames from FakeAVCeleb [5] dataset with open mouths. Here f represents frame number. The first three columns present consecutive frames for local comparison, while the last two columns offer a broader perspective by displaying frames with similar poses from the entire video, defined as global inconsistencies in our paper. The deepfakes exhibit more pronounced inconsistencies in aspects such as mouth shape, coloration, dental structure, and tongue appearance.
@@
Deepfakes are broadly categorized into two types based on the regions of manipulation: entire-face synthesis and partial manipulation. Entire-face synthesis encompasses techniques such as face-swapping and talking head generation. On the other hand, partial manipulation primarily involves lip-syncing videos. In these videos, the lip movements of an individual are altered to match a specific audio track, creating the illusion that they are speaking words they originally did not. Lip-synced deepfakes focus on modifying only a person’s mouth area, which facilitates the creation of high-quality deepfakes. A recent investigation [6] indicates that lip-synced deepfakes pose a greater challenge for human detection than face-swapping deepfakes. However, unlike the extensively researched face-swap deepfake detection [7, 8, 9, 10, 11], identifying lip-syncing deepfakes presents greater challenges and is less explored.
Figure 2: End to End pipeline of the proposed LIPINC model.
@@
II Related Work
II-A DeepFake Detection
Most deepfake detection approaches focus on entire-face synthesis in videos, identifying visual anomalies like face warping [9], facial movements [16], discrepancies between face and context [17], and inconsistencies in face identity [18]. These methods extract distinctive features from the entire face and are less effective when dealing with partially manipulated lip-syncing videos. For lip-sync deepfake detection, various techniques are employed to exploit audio-visual mismatches. Shahzad et al.[19] proposed a method to detect fabricated videos by contrasting the real lip sequence from a video with a synthetic lip sequence created based on audio using a lip-sync model [3]. Feng et al.[13] investigated audio-visual signal inconsistencies in forged videos through anomaly detection, but this approach might miss manipulations that achieve good synchronization [13]. Alternatively, Haliassos et al.[14] identified forged videos by detecting semantic anomalies in mouth movements from video segments, focusing on irregularities in mouth motion from lower face regions, which may be less effective in videos with minimal mouth movement. LipForensics [14], in contrast, learned mouth motion patterns using lipreading networks from lower-face clips for deepfake identification. Tolosana et al.[15] examined visual artifacts across various facial regions (eyes, nose, and mouth), utilizing spatial indicators to differentiate between real and fake videos.
II-B Lip-syncing Deepfake Datasets
Early deepfake datasets are for entire-face synthesis, such as face-swapping and face-reenactment [2, 20]. The emergence of Wav2Lip [3] in 2020 marked a significant advancement in lip-syncing deepfakes, and led to the development of several datasets aimed at detecting partially manipulated lip-syncing videos. For instance, the FakeAVCeleb [5] dataset is a comprehensive audio-video deepfake detection resource featuring accurately lip-synced videos. This dataset encompasses three categories of audio-video deepfakes sourced from English-language YouTube videos, including lip-syncing [3], face-swapping [1, 21], and voice cloning [22]. Additionally, the KODF [23] dataset presents a vast collection of both generated and authentic videos with Korean subjects, with the fake videos produced using Wav2Lip and various face-swapping models. The CMDFD [24] dataset covers two lip-syncing models and two talking head generation models for cross-modal Deepfake detection. Note that these existing datasets encompass only a subset of lip-syncing videos created with a single model.
III Method
@@
The overall pipeline of the proposed LIPINC model is depicted in Fig. 2. It comprises two primary modules: 1) Local and Global Mouth Frame Extractor, which isolates adjacent and similarly posed mouth segments based on mouth openness throughout the video sequence; and 2) Mouth Spatial-Temporal Inconsistency Extractor (MSTIE), responsible for encoding and learning distinctive inconsistency features within and across frames, focusing on both color and structure aspects. The output of the MSTIE module is then directed to a binary classifier to determine the probabilities of each class.
III-A Local and Global Mouth Frame Extractor
Given an input video, we first use a face detector, such as Dlib [25], to crop and align the face in each video. Then we use the facial landmarks to extract the mouth region, as shown in Fig. 3. We propose to extract multiple frames with open mouths since they contain more inconsistencies regarding deepfakes. For the adjacent mouth extraction as local inconsistency analysis, we obtain a sequence of $L$ adjacent frames with open mouth, which are selected based on a large distance between landmarks 63 and 67 in Fig. 3. For the global inconsistency representation, we search for $G$ similar pose frames identical to the mouth openness in local frames from the rest of the video. The similar pose frames are extracted by comparing the height and width of opened-lips with similar face orientations. To ensure a global matching and prevent selected frames from being adjacent, we set the extraction of similar-pose frames to have a minimum time gap of 0.09 seconds between them. We then concatenate the adjacent and similar-pose frames to form our color sequence C with $N$ frames where $\textbf{C}\in R^{(N*H*W*3)}$, 3 is the channel number, and $N=L+G$. To emphasize the consistency description at the mouth structure level, we further add the structure branch with residual sequence S, where $\textbf{S}\in R^{((N-1)*H*W*3)}$ is calculated by
$S_{t}=C_{t+1}-C_{t}\;(1\leq t<N)$ (1)
The structural branch is dedicated to capturing the inconsistencies in mouth and teeth shapes across different frames. The color and structure sequences with local and global mouth frames are fed into the following inconsistency feature extraction module.
Figure 3: Mouth region landmarks detected by Dlib [25]. Orange colors denote the landmarks for mouth openness measurement and matching.
@@ III-A Local and Global Mouth Frame Extractor
TABLE I: In-domain testing on FakeAVCeleb dataset [5] and cross-domain testing on KODF [23] and LSR+W2L. We compare our model with state-of-the-art detectors regarding Precision, Accuracy, Average Precision and AUC score. Here RI means Re-Implemented. Note that the results sourced from existing methods used much more training data than us for in-domain cross-type testing.
III-B Mouth Spatial-Temporal Inconsistency Extractor
This MSTIE module is designed to encode the color and structure frames and learn discriminative features representing spatial and temporal inconsistency levels for deepfake detection. We use the 3D-CNN [28] model to produce the spatial-temporal features $\textbf{v}_{ci}$ and $\textbf{v}_{si}$ for the color and the structure sequence, respectively, where $i$ is the number of input video. These vectors are then passed through a cross-attention [29] module which interconnects the color and structure branches such that $K_{ci}=V_{ci}=\textbf{v}_{ci},$ $Q_{ci}=\textbf{v}_{si}$ and $K_{si}=V_{si}=\textbf{v}_{si}$, $Q_{si}=\textbf{v}_{ci}$, where $K,V$ and $Q$ represent the keys, values, and queries. The cross attention for color $A_{c}$ and structure $A_{s}$ branch is calculated as follows:
$A_{ci}=\text{softmax}\left(\frac{Q_{ci}K_{ci}^{T}}{\sqrt{d_{qc}}}\right)V_{ci}$ (2)
$A_{si}=\text{softmax}\left(\frac{Q_{si}K_{si}^{T}}{\sqrt{d_{qs}}}\right)V_{si}$ (3)
@@ III-B Mouth Spatial-Temporal Inconsistency Extractor
$A_{fi}=\text{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{Q}}}\right)V_{i}+\textbf{v}_{si}$ (4)
@@
III-C Loss Function
@@
$L_{IL}=-\frac{1}{M}\sum_{i=1}^{M}y_{i}log(AvgS_{i})+(1-y_{i})log(1-AvgS_{i})$ (5)
@@
IV Experiments
Figure 4: Ablation analysis on LSR+W2L and KODF-LS Dataset based on AUC scores. Here L, G, C, and S refer to Local, Global, Color, and Structure frames, respectively. IL is Inconsistency Loss.
IV-A Settings
Datasets. We use the FakeAVCeleb dataset[5] to train and validate our model. The FakeAVCeleb dataset consists of 500 real unique frontal face videos and more than 20,000 multi-modal deepfake videos. We split the dataset into two categories: Lip-synced (FakeAV-LS) deepfakes generated by the Wav2lip [3] model and Face-swap (FakeAV-FS) deepfakes that have not used Wav2lip. Since our method mainly focuses on lip-syncing deepfakes detection, we have divided the lip-synced deepfakes in this dataset into a non-overlapped training set and a test set. We use two more datasets to evaluate our model’s generalization ability. The first one is the KODF dataset [23], which includes 62,166 real videos and 175,776 deepfake videos. Similar to the FakeAVCeleb dataset, deepfake videos contain lip-syncing and entire-face synthesis. We created 3 test sets from the dataset based on the deepfake generation method. 1) KODF-LS: audio-driven lip-synced deepfakes from [3, 31]; 2) KODF-FSGAN: Faceswap deepfakes generated using [21]; 3) KODF-DFL:Faceswap deepfakes generated using [32]. We randomly selected a subset of real and fake videos for evaluation following previous studies [13], and use 800 samples for each class. Secondly, we have generated a new dataset called LSR+W2L, which is specifically tailored for lip-syncing detection. We consider two models, Wav2Lip and Wav2Lip with GAN model [3]. This dataset contains 1,000 real videos from the LRS2 dataset [33] and 1,000 fake videos.
Evaluation Metrics. The proposed LIPINC model is evaluated using four common metrics in deepfake detection, including Precision, Accuracy, Average Precision (AP), and Area Under the Curve (AUC) scores [34].
@@ IV-B Comparison under In-domain Testing
In Table I, we first evaluate our model on the FakeAVCeleb testing set (FakeAV-LS) [5] and compared its performance with the state-of-the-Art (SOTA) methods. The proposed LIPINC model achieves the best performance in terms of accuracy, precision and is tied with the FTCN [7] model for the best performance in terms of AUC scores. Our model outperforms the best SOTA method Intra-Cross-modal [26] in terms of AUC scores by 1% and performs slightly worse than it in AP. Overall, the LIPINC model demonstrates outstanding performance in detecting lip-syncing videos under in-domain testing, thanks to the utilization of spatial-temporal mouth inconsistency features.
IV-C Comparison under Cross-domain Testing
@@
We further evaluate our model on KODF-LS[23] and LSR+W2L datasets to evaluate the generalization capability of our model. The results are shown in Table I. Note that our model was trained on the FakeAVCeleb training set (FakeAV-LS), while the SOTA results cited from existing papers were trained on the whole FakeAVCeleb dataset [5] as introduced in [13]. On our newly created LSR+W2L dataset, the proposed LIPINC model outperforms all the baselines in terms of AP and AUC scores. The RealForensics model achieves better performance than our method on the KODF dataset due to its utilization of a large training dataset. Our method achieves an AUC of over 87% on both datasets, demonstrating strong generalization in detecting unknown lip-syncing videos.
IV-D Cross-type Deepfake Testing